OpenStudy (anonymous):

Tough Calc Question: Let r>0. The equations below parametrize an astroid x= r(cos^3)t and y= r(sin^3)t for 0<=t<=2pi a. find the length of the asteroid b. show that the astroid is alternatively the graph of x^(2/3) + y^(2/3) = r^(2/3) c. Find the area S of the surface generated by revolving C about the x axis. Been stuck on this problem for a while.. any help would be appreciated! Thanks.

OpenStudy (sirm3d):

(A) \[\text{length of curve}=\Large \int_0^{2\pi}\sqrt{\left(\frac{dx}{dt}\right)^2+\left(\frac{dy}{dt}\right)^2}\;\mathrm dt\]

OpenStudy (sirm3d):

(B) \[\Large {x=r\cos^3t\\x^{1/3}=r^{1/3}\cos t\\x^{2/3}=r^{2/3}\cos^2 t\\}\] do the same on \(y\), and use the trigonometric identity \[\Large \cos^2 t + \sin^2 t=1\]

OpenStudy (anonymous):

C \[S = \int\limits_{0}^{2 \pi} 2 \pi y \sqrt{\left(\frac{dx}{dt}\right)^2 + \left(\frac{dy}{dt}\right)^2}\:dt\]